# TimeCapsule: Solving the Jigsaw Puzzle of Long-Term Time Series Forecasting with Compressed Predictive Representations

### v1

@@ 5.3. Efficiency study
Despite TimeCapsule containing multiple transformer blocks along three dimensions, it maintains a reasonable computational complexity due to the use of dimensionality compression and MLPs. To confirm this, we conduct efficiency comparisons on two datasets, Weather and Traffic, analyzing both FLOPs and parameter counts. As shown in Fig. 5, our model demonstrates clear advantages in computation speed, though memory usage can sometimes be high.
Nevertheless, a principal advantage of TimeCapsule is the performance-efficient balance attained by its asymmetric structure, which permits a substantial reduction in parameters without much sacrifice in performance. For instance, we have reduced the hyperparameters, including the number of layers and the number of neurons, for the purpose of forecasting on the Traffic dataset. As shown in the figure, a light version of TimeCapsule attains an MSE of 0.373, compared to the original result of 0.361 (the version incorporating JEPA training). However, in the meantime, both FLOPs and memory usage undergo substantial reduction.

### v3

@@ 5.3. Efficiency study
Despite TimeCapsule containing multiple transformer blocks in three dimensions, it maintains a reasonable computational complexity due to the use of dimensionality compression and MLPs. To confirm this, we conduct efficiency comparisons on two datasets, Weather and Traffic, analyzing both FLOPs and parameter counts. As shown in Fig. 7, our model demonstrates clear advantages in computation speed, though memory usage can sometimes be high.
Nevertheless, a principal advantage of TimeCapsule is the efficient-performance balance attained by its asymmetric structure, which allows a substantial reduction in parameters without much sacrifice in performance. For instance, we have reduced the hyperparameters, including the number of layers and the number of neurons, for the purpose of forecasting on the Traffic dataset. As shown in the figure, a light version of TimeCapsule attains an MSE of 0.373, compared to the original result of 0.361 (the version incorporating JEPA training). However, in the meantime, both FLOPs and memory usage undergo substantial reduction.